OpenStudy (brandyscott29):

How many solutions does the following system have? -3x-2y=-12 9x+6y=-6 I know that they have no solutions but I just don't understand why.. Help?

OpenStudy (mhchen):

You can make the 2 equations the same on one side. If you multiply the top equation by -3, It becomes: -3(-3x-2y) = (-12)-3 9x + 6y = 36 ^^^^^^ You see how we got 9x + 6y = something? In the other equation, we ALSO got 9x + 6y = something. So that means the 'something' has to be the same, but -6 does NOT equal to 36. Therefore this system of equations is impossible to solve.